Calculate the magnitude of the electric field at one corner of a square 2.12 m on a side if the other three corners are occupied by 4.25x10-6 C charges. Express your answer to three significant figures and include the appropriate units E-Value Units Submit Request Answer ▼ Part B What is the direction of the electric field at the corner? along the line between the corner and the center of the square toward the center along the side of the square between the corner and one of the charges outward of the charge O along the side of the square between the corner and one of the charges toward the charge along the line between the corner and the center of the square outward of the center Submit Request Answer
Answer #1

Solution)

Part a) F1= F3= kq/a^2 =(9*10^9)*(4.25*10^-6)/2.12^2

= 8510.59 N/C

F2= kq/2a^2 =4255.59 N/C

Now, Fr= sqrt(F1^2 + F3^3) + F2

Fr= 16291.38 N/C

=======

Part b)

Its along the line between the corner and the center of the square outward of the center.

=====